OpenStudy (miracle223):

can any body tell me the answer? Catherine paid $218.00 for an MP3 player. If the price paid includes a 9% sales tax, which of the following equations can be used to determine the price of the MP3 Player before tax?

OpenStudy (miracle223):

(Let x represent the cost of the MP3 player and y represent the total cost after tax.) y = 0.91x y = 1.9x y = x + 9x y = 1.09x these are the choices

Directrix (directrix):

x represents the cost of the MP3 player y represents the total cost after tax 9% = .09 x = 1*x ---------------------------- The cost of the player + .09 * cost of the player = 218 y = x + .09*x y = 1*x + .09*x y = ? @miracle223